OpenStudy (anonymous):

A round trip flight from Seattle, Washington to Orlando, Florida is one of the longest flights in the continental United States. Due to the jet stream that blows from west to east, a plane flying nonstop from Seattle to Orlando flies at an average speed of 523 miles per hour. The same plane flying from Orlando to Seattle travels at an average of 415 miles per hour. The full round trip takes exactly 11 hours. How far is Seattle from Orlando, to the nearest mile?

OpenStudy (anonymous):

How far is Seattle from Orlando? Let d = one way distance (?), t = total trip time (11 hours), So = speed out (523 mph), Sb = speed back (415 mph) time out = d/So = d/523 time back = d/Sb = d/415 t = time out + time back 11 = d/523 + d/415 find common denominator here 11 = (415d + 523d)/(523*415) 11*523*415 = 415d + 523d 2387495 = 938d d = 2545 miles

OpenStudy (anonymous):

Thank you very much. Your answer is correct, I checked in the textbook answers. I just needed a systematic way of solving it. :)

OpenStudy (anonymous):

Did I get an medal?

OpenStudy (mertsj):

Be altruistic

OpenStudy (anonymous):

Yeah you got a medal.

OpenStudy (radar):

Here is a medal for you kumar2006, I liked the layout of your solution and of course the correct answer.